OpenStudy (anonymous):

What are the first five terms in the arithmetic sequence An=6n, starting with n=1? I am thinking it is 7,8,9,10,11, but I am not sure

OpenStudy (amistre64):

what is 6n, when n=1? what is 6n, when n=2? what is 6n, when n=3? ... what is 6n, when n=5?

OpenStudy (anonymous):

OH!! I see! Not as hard as I thought! I have never seen this before tho. Thank you!

OpenStudy (amistre64):

i recall the first time i seen it. it does take a little bit of getting used to :)

OpenStudy (anonymous):

ok so if I have the first five terms in the geometric sequence An=(2)^n-1 what the heck does that mean?

OpenStudy (amistre64):

all your doing is plugging n=1,2,3,4,5 into the spot reserved for the n, and evaluating

OpenStudy (anonymous):

so it would be 2, 4, 8, 16, 64?

OpenStudy (amistre64):

close; hard to read the ascii notation \[2^{n-1}\] or \[2^n-1\] ???

OpenStudy (anonymous):

\[2^{n-1?} when n=1\]

OpenStudy (anonymous):

when n=1

OpenStudy (amistre64):

then the output is simply \[2^{n-1}=2^0,2^1,2^2,2^3,2^4\]

OpenStudy (anonymous):

ah I get it! Thank you! Boy that just sounded a lot harder than it was

OpenStudy (amistre64):

if you are better used to playing with functions of x, instead of n, the name of the variable doesnt matter. \[f(x)=2^{x-1}~;~x=1,2,3,4,5\]is the same thing

OpenStudy (anonymous):

\[\sum_{k}^{5?}=then \it has a little 1 then a k \in the \middle and a little 2 \above that and I have never seen that symbol either\]

OpenStudy (amistre64):

\[\sum_{k}^{5?}=\]then \it has a little 1 then a k \in the \middle and a little 2 \above that and I have never seen that symbol either

OpenStudy (amistre64):

might be better to use the draw feature if the equation editor is giving you issues :)

OpenStudy (amistre64):

the symbol "\(\sum\)" is just a greek capital letter "S", and represents the SUM of a stated rule or expression.

OpenStudy (anonymous):